For the given rational function f, find all values of x for which f(x) has the indicated value.
f(x) = 5x + 21/x ; f(x) = -38

Answer by london maths tutor(243)

f(x) = 5x + 21/x ; f(x) = -38
-38 = 5x + (21/x)
-38x = 5x^2 + 21
5x^2 + 38x + 21 = 0
(5x + 3)(x + 7) = 0
Therefore,
5x+3 = 0
x = -3/5
or
x+7 = 0
x=-7
Answer: x = -3/5 and -7